OpenStudy (anonymous):

m varies jointly with the square of n and the square root of q. If m = 252 when n = 2 and q = 49, find m when n = 7 and q = 4.

OpenStudy (amistre64):

m=n^2 sqrt(q) k 252 --------- = k 7^2 sqrt(4)

OpenStudy (amistre64):

k = 252/49.2 = 252/98 m=(252/98) 7^2....read the prob bacwards lol

OpenStudy (amistre64):

252 = k.2^2.sqrt(49) 252 = k.4.7 k = 252/28

OpenStudy (amistre64):

m = (252/28).7^2.sqrt(4) m = 252(49)(2) --------- 7(4)

OpenStudy (amistre64):

m = 252(7)/2 = 126(7) = 700+140+42 m = 882